[figure: 9.7 cm]

SPIN REORIENTATION AND THERMAL EXPANSION OF

LATTICE PARAMETERS IN TmFeO_{3}

A. Bombik^{a}, J. Kusz^{b}, A.W. Pacyna^{c}, J. Przewoźnik^{a}

a Faculty of Phys. Nucl. Techniques, Mining and Metallurgy University, Kraków

b Institute of Physics, University of Silesia, Katowice, Poland c Henryk Niewodniczański Institute of Nuclear Physics, Kraków, Poland

Temperature dependences of lattice parameters were examined for both single crystal and powder samples of TmFeO_{3} orthoferrite. The single crystal study was performed on X-ray four circles Huber diffractometer between 10 K and 300 K while powder measurements were carried out using a Simens D-5000 diffractometer from 70 K up to 300 K. Lattice constant temperature dependences determined from single crystal and powder measurements are consistent in the covered temperature range. The thermal expansion coefficients are anisotropic. For the b direction it is almost constant at all measured temperatures while for a and c ones they show considerable temperature dependences. Expansion coefficients are positive above spin reorientation region and negative below. Lattice parameters a and c reach their minimum values within spin reorientation range (approx. 80-100 K). The low temperature behaviour of a and c lattice parameters seems to originate from a magnetostriction effect of TmFeO_{3}.

This work was supported by research grant No 2 P03B 035 14 from State Committee for Scientific Research.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Andrzej W. Pacyna

Address for correspondence : Instytut Fizyki Ja̧drowej im. Henryka Niewodniczańskiego ul. Walerego Eljasza Radzikowskiego 152 31-342 Kraków

Email address : Andrzej.Pacyna@ifj.edu.pl